Shift report 20011125 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Nov 25 2001 - 16:01:17 EST

  • Next message: Flemming Videbaek: "End of Au-Au run"

    ------------------------------------------------------------
    Date: Sun Nov 25 16:01:17 2001, Shift: 08:00-16:00
    Supervisor: rw, with: fv,mm
    Log book#11 pages 33 - 38
    
    Shift Summary:
    ------------------------------------------------------------
    1.22Gev beam was ready at 10:00. FV and MM worked for ZDC and
    trigger in the morning. Physics run started at 11:00. Beam stored 
    and, after 10min., cogged (without energy ramping), waiting for
    being ready for voltage setup for all four experiments and lasted
    about 1h for data taking and dumped. In 10min. next beam was stored
    and cogged. This cycle was repeated beautifully without break.
    Hope to see this beautiful cycle for 100Gev next year. ZDC rate was 
    about 15 in the beggining.
    
    2. 	run 	trigger1
    	5992    10.7K		TPM1,TPM2,T1,T2 ON 
    				FS 12deg-1/10A
    				MRS 90deg-2A
    	5993	13.3K
    	5995	11.4K
    	5996	22K
    
    3. no new detecotr problems
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Nov 25 2001 - 16:01:50 EST